1.A jury of 12 people is to be randomly selected from a group of 30 people (18 women and 12 men).  Let the random variable X = “number of women on the jury of 12.”  Find (using R):

      

      a) P(X = 6)

      b) P(X ≤ 8)

      c) P(5 ≤ X ≤ 10)

      d) P(no men on the jury)
Answer #1

Since , X =“number of women on the jury of 12.” follows Binomial distribution ie Bin ( n,p)

where n = 12 and p = 18/30 ie p= 0.6. X is Binomial since there can be only 2 outcomes of an event ( women selected or not selected). Also, the trials are independent of each other.

Attached please find the R Code :

RStudia Eile Edit Code V ls Session uild Deug rofil Topls Help * 옛 甘▼1日申1曲1-Go to file/function | | 63 -Addins Project (None) q1.Rx Environment History Connections Source on Save a 目 → Run-+ Source ▼--で日→Import Dataset List # x-number of wome n on the jury of 12 Global Environment 2 n<- 12 3 p<-0.6 4 5 a <- dbinom(6, n, p) 6 print (a) 7 values # (a) (X=6) 0.176579149824 0.77466271744 0.923099037696 0.002176782336 12 0.6 P # b (b) <-pbi nom (8, n , p , 1 over. tail P (X<-8) = TRUE) 9 print (b) 10 # (c) P(5 x 10) 11 c<- pbinom 10,n,p, lower.tail TRUE)-pbinom (4,n,p,lower.tail TRUE) 12 print (c) 13 14 # (d) P(no men on the jury) -P(X=12) 15 d <-dbinom(12,n,p) 16 print (d) 17 17:1 ( op Level R Script 수 Console Terminal Files Plots Packages Help Viewer > print(a) qbinom [1] 0.1765791 R: I he Binomial Distribution ais Find in Topic > print(b) [1] 0.7746627 > print (c) [1] 0.923099 > print (d) [1] 0.002176782 3 rob, log = FALSE) rob, lower.tailTRUF, log.pFATSE) 1156 PM Type here to scarch